A random variable follows a binomial distribution with a probability of success equal to 0.72. For a sample size of n=10​, find the values below.
a. the probability of exactly 5 successes
b. the probability of 6 or more successes
c. the probability of exactly 9 successes
d. the expected value of the random variable

Answer by ikleyn(52754)

.

Go to web-site https://stattrek.com/online-calculator/binomial.aspx

and use free of charge online calculator there, intended for common use.


This calculator has nice, simple and convenient interface, which serves for all listed options
in your post, so even a beginner student can easily solve this assignment in parts  a),  b)  and  c).


With this calculator, solving problems like  a),  b)  and  c) is only a matter of pressing buttons.


Regarding part  d),  use the formula for binomial distribution

        expected value = n*p = 10*0.72 = 7.2.         ANSWER